Year Cash inflow (R)

3 45 000

6 90 000

9 115 000

The applicable interest rate is 11,59% per year. The present value of  the cash outflows is R95 000

The future value of the cash inflows is approximately

[1] R326 950. [2] R271 470. [3] R169 330. [4] R218 000. [5] R250 000


1
Expert's answer
2021-07-22T10:01:26-0400

In this case, we need to compute the future value of cash inflows at the end of year 9. Thus, the present value of cash outflow is irrelevant. 

Calculating the future value of cash inflows:

"Future value=\frac{ C3}{(1+r)^6}+\frac{C6}{(1+r)^3}+\frac{C9}{(1+r)^0} \\=\frac{45,000}{(1+0.1159)^6}+\frac{90,000}{(1+0.1159)^3}+\frac{115,000}{(1+0.1159)^0}\\=326,950"

Where:

Cash inflow in Year 3 (C3)=45,000,

Cash inflow in Year 6 (C6)=90,000, 

Cash inflow in Year 9 (C9) = 115,000, and

Interest rate (r) = 0.1159 or 11.59%, 

 

Thus, option (1) is correct.